# Cognitive Paradigms for Evaluating VLMs on Visual Reasoning Task

### v1

@@ B.1 Bongard OpenWorld Dataset
We utilize a subset of 500 cases from the Bongard OpenWorld dataset Wu et al. (2024), a benchmark for evaluating few-shot reasoning in machine vision. Each case presents a set of positive and negative examples, distinguished by an underlying "commonsense" rule. These rules, represented by numerical values (0-9), encode different visual and spatial concepts, as detailed below.
B.2 Commonsense Value Categories
@@
B.3 Commonsense Value Distribution in Our Subset
Table A.5 provides a detailed breakdown of the distribution of commonsense values within our 500 test-cases subset. As observed, the value ’0’ ("Anything else," marked as commonsense) is significantly more prevalent than other values. This distribution should be considered when analyzing the model’s performance, as it may indicate a bias towards learning certain types of rules, particularly those related to the "Anything else" category.
@@ B.3 Commonsense Value Distribution in Our Subset
Table A.4: Distribution of Commonsense ID in Bongard OpenWorld Dataset
B.4 Dataset Availability
@@
Appendix C Evaluation Metrics
@@
Appendix D Model Details
D.1 Vision-Language Models (VLMs)
  • ChatGPT (GPT-4o with vision): OpenAI’s GPT-4o, accessed with vision capabilities via the openai library, processes both images and text seamlessly. It OpenAI (2024) excels in reasoning and instruction-following tasks, offering robust text generation and broad general knowledge. However, its closed-source nature and cost make it less accessible for extensive experimentation.
  • Gemini (2.0-flash-exp): Google’s Gemini is a multimodal model Google (2024) emphasizing speed and computational efficiency in its "flash" version, accessed through the google.generativeai library. While it provides rapid inference capabilities, it is an experimental model with limited public documentation, restricting detailed architectural insights.
D.2 Model Configurations and Hyperparameters
For closed-source models, we utilized API calls by installing the openai, mistralai (for Pixtral), and google.generativeai libraries, enabling access to ChatGPT (GPT-4o with vision) OpenAI (2024), Pixtral (12B) Agrawal et al. (2024), and Gemini (2.0-flash-exp) Google (2024), respectively. The open-source models, on the other hand, were accessed using the ollama Ollama (2025) library.
  • Input Format: Bongard problems were presented by encoding images in base64 and embedding them within text prompts. For instance: "Here are two sets of images. The first set represents positive examples, and the second set represents negative examples of a concept. What is the rule that distinguishes the two sets? "
  • Image Size: For OpenAI and Gemini models, the default image sizes provided by their APIs were used. For all other models, the maximum image size was restricted to 1024 pixels to ensure uniformity.
Appendix E Results
Concept Category ID GPT 4.0 Gemini 2.0
Anything else 0 92.88 94.23
Human-Object Interaction (HOI) 1 86.67 92.86
Taste / Nutrition / Food 2 100.00 85.71
Color / Material / Shape 3 88.89 91.67
Functionality / Status / Affordance 4 100.00 100.00
And / Or / Not 5 90.00 80.00
Factual Knowledge 6 90.00 90.00
Meta Class 7 100.00 100.00
Relationship 8 100.00 100.00
Unusual Observations 9 92.31 92.31
Table A.5: Accuracy per Commonsense ID and Categories in Bongard OpenWorld
Gemini 2.0 GPT-4o
Category Mean Std Mean Std
Positive 0.915 0.02 0.9022 0.02
Negative 0.868 0.02 0.8659 0.02
Table A.6: Semantic Similarity between query image Descriptions and Identified Rules
SL NO Task ID Caption Reason
1 0021_neg_0 Cars on the city streets at night The image is synthetic, not a real-world photograph.
2 0014_neg_0 A person playing a guitar. The image is synthetic, lacking real-world context.
3 0033_neg_0 A bicycle is placed in the corner The image is synthetic and does not represent real-world scenarios.
4 0037_neg_0 The girl has long and thin braids on her head. While technically correct, the braids are incorrectly interpreted or misrepresented.
5 0076_pos_0 Various kinds of rings The classification error arises due to confusion between antique and artistic presentation styles.
6 0076_neg_0 Various kinds of rings The classification error arises due to confusion between antique and artistic presentation styles.
7 0082_neg_0 Live coral on the sea floor. The analysis is too fine-grained.
8 0084_neg_0 A wooden fence surrounding a grassy field. Failure to accurately identify the wooden fence in the image.
9 0112_neg_0 A wooden floor in the living room. The analysis is too fine-grained, and the image does not match the caption.
10 0117_neg_0 Colorful ribbons. While ribbons are present, they are not colorful as described.
11 0122_neg_0 A satellite view of Earth. Unable to recognize or correctly interpret the satellite image.
12 0136_pos_0 Spectator seats view in the stadium. The analysis is overly detailed, leading to a categorization error.
13 0213_neg_0 Checkerboard pattern fabrics Only the checkerboard pattern was identified, missing other key aspects.
14 0234_neg_0 A beautiful stone sculpture The image is misclassified as a stone obelisk monument rather than a sculpture.
15 0247_pos_0 Small river filled with reeds Although the identification is correct, the error stems from overly fine-grained analysis.
Table A.7: Task IDs and descriptions where GPT-4o and Gemini both incorrectly categorize.
E.1 Holistic Analysis
@@
E.2 Deductive Rule Learning
First-Stage Prompt (Rule Extraction)
@@ E.2 Deductive Rule Learning
Second-Stage Prompt (Rule Application)
@@
E.3 Componential Analysis
First-Stage Prompt (Image Description Generation)
@@ E.3 Componential Analysis
Second-Stage Prompt (Rule Derivation Instruction)
def system_eval_prompt():
return """
You are an expert in visual reasoning and analysis, working with complex datasets such as the Bongard dataset. Your task is to analyze descriptions of images from two categories (cat_2 and cat_1) along with a query image, provided in JSON format.
The goal is to:
1. Identify the **common characteristics or patterns** in the positive samples (cat_2) that distinctly separate them from the negative samples (cat_1).
2. Derive a **clear and concise rule** that defines the positive samples.
3. Apply this rule to the query image and categorize it as belonging to either cat_1 or cat_2.
Ensure the output is clear, concise, and follows the provided format:
- **Analysis**: Provide a brief analysis of the distinguishing characteristics or patterns.
- **Rule**: State the rule that separates cat_2 from cat_1.
- **Query Image**: Analyze the query image based on the derived rule.
- **Conclusion**: Clearly state whether the query image belongs to cat_1 or cat_2.
Respond in the format provided, and avoid unnecessary explanations or text.
Third-Stage Prompt (Rule Application and Classification)
@@
The image descriptions for the positive samples, negative samples, and the query image are provided in JSON format. Analyze the common patterns or characteristics in the cat_2 samples that distinguish them from cat_1 samples.
@@
2. Apply this rule to categorize the query image.
@@
- **Query Image**: (Query image details)

### v2

@@ A.3.1 Bongard OpenWorld Dataset
We utilize a subset of 500 test cases from the Bongard OpenWorld dataset Wu et al. (2024), a benchmark for evaluating few-shot reasoning in machine vision. Each of the 1001 Bongard samples in this dataset contains 7 positive and 7 negative images, distinguished by an underlying "commonsense" rule represented by numerical values (0-9). To create our evaluation set, we followed a specific sampling strategy. Each test case in our subset consists of 6 positive images, 6 negative images, and 1 query image. From each of the original 1001 Bongard samples, we can generate up to 14 unique, balanced test cases using combinations. For our evaluation, we selected the first 250 Bongard samples from the full dataset. From each of these 250 samples, we selected one positive and one negative test case, resulting in a total of 500 test cases for our analysis. The specific samples used in our analysis will be made publicly available upon acceptance of this work. These rules encode different visual and spatial concepts, as detailed below.
Figure A.1: Illustration of a Bongard OpenWorld dataset showing three positive and negative samples along with a query image. Distinguishing rule in this example is A group photo at a wedding reception. Based on the rule, query image belongs to negative category.
A.3.2 Commonsense Value Categories
@@
A.3.3 Commonsense Value Distribution in Our Subset
Table A.2 provides a detailed breakdown of the distribution of commonsense values within our 500 test-cases subset. As observed, the value ‘0’ (‘Anything else’, marked as commonsense) is significantly more prevalent than other values. This distribution should be considered when analyzing the model’s performance, as it may indicate a bias towards learning certain types of rules, particularly those related to the ‘Anything else’ category.
@@ A.3.3 Commonsense Value Distribution in Our Subset
Table A.2: Distribution of Commonsense ID in Bongard OpenWorld Dataset test cases used for experiments
A.3.4 Dataset Availability
@@
Appendix A.4 Model Details
A.4.1 Vision-Language Models (VLMs)
  • ChatGPT (GPT-4o with vision): OpenAI’s GPT-4o, accessed with vision capabilities via the openai library, processes both images and text seamlessly. It OpenAI (2024) excels in reasoning and instruction-following tasks, offering robust text generation and broad general knowledge. However, its closed-source nature and cost make it less accessible for extensive experimentation. Accessed version of the model is 2024-08-06.
  • Gemini (2.0-flash-exp): Google’s Gemini is a multimodal model Google (2024) emphasizing speed and computational efficiency in its “flash” version, accessed through the google.generativeai library. While it provides rapid inference capabilities, it is an experimental model with limited public documentation, restricting detailed architectural insights.
A.4.2 Model Configurations and Hyperparameters
For closed-source models, we utilized API calls by installing the openai, mistralai (for Pixtral), and google.generativeai libraries, enabling access to ChatGPT (GPT-4o with vision) OpenAI (2024), Pixtral (12B) Agrawal et al. (2024), and Gemini 2.0 (flash-exp) Google (2024), respectively. The open-source models, on the other hand, were accessed using the ollama Ollama (2025) library.
  • Input Format: Bongard problems were presented by encoding images in base64 and embedding them within text prompts. For instance: “Here are two sets of images. The first set represents positive examples, and the second set represents negative examples of a concept. What is the rule that distinguishes the two sets? ”
  • Image Size: For OpenAI and Gemini 2.0 models, the default image sizes provided by their APIs were used. For all other models, the maximum image size was restricted to 1024 pixels to ensure uniformity.
A.4.3 Evaluation Metrics
@@
Appendix A.5 Model Prompts
A.5.1 Direct Visual Rule Learning
The prompt used for the Direct Visual Rule Learning paradigm is designed to elicit a holistic analysis of the provided images, encouraging the model to identify a distinguishing rule and apply it to the query image. The prompt emphasizes the distinction between positive ($cat\_2$) and negative ($cat\_1$) examples and guides the model to provide a structured output containing its analysis, the identified rule, details about the query image, and the final classification.
@@ A.5.1 Direct Visual Rule Learning
A.5.2 Deductive Rule Learning
The Deductive Rule Learning paradigm employs a two-stage prompting strategy. The first stage focuses on rule extraction from positive and negative examples, while the second stage applies the extracted rule to classify a query image. The prompts for each stage are detailed below.
A.5.2.1 First-Stage Prompt (Rule Extraction)
This prompt guides the model to identify and summarize a distinguishing rule based on provided positive and negative examples. It emphasizes conciseness in the rule summary.
@@ A.5.2.1 First-Stage Prompt (Rule Extraction)
A.5.2.2 Second-Stage Prompt (Rule Application)
This prompt presents the previously extracted rule summary and a query image, prompting the model to classify the image based on the rule. It reinforces the Bongard problem context and requests a structured output.
@@ A.5.2.2 Second-Stage Prompt (Rule Application)
A.5.3 Componential Analysis
The Componential Analysis paradigm also uses a two-stage prompting strategy. The first stage generates detailed image descriptions, while the second stage derives a rule from these descriptions and applies it to a query image. The specific prompts for each stage are presented below.
A.5.3.1 First-Stage Prompt (Image Description Generation)
This prompt instructs the model to generate a comprehensive, hierarchical description of a given image in JSON format. It guides the model to cover various aspects of the image, from scene and objects to activities and contextual elements, facilitating detailed comparative analysis in the subsequent stage.
@@ A.5.3.1 First-Stage Prompt (Image Description Generation)
A.5.3.2 Second-Stage Prompt (Rule Derivation Instruction)
This prompt guides the model to analyze the JSON descriptions generated in the first stage, derive a distinguishing rule, and apply it to classify a query image. It emphasizes the use of the provided JSON format and requests a structured output.
@@ A.5.3.2 Second-Stage Prompt (Rule Derivation Instruction)
The image descriptions for the positive samples, negative samples, and the test image are provided in JSON format. Analyze the common patterns or characteristics in the cat_2 samples that distinguish them from cat_1 samples.
@@
2. Apply this rule to categorize the test image.
@@
- **Test Image**: (Test image details)
@@
Appendix A.6 Results
This appendix provides supplementary information and detailed results for the experiments described in the main paper. We present additional performance breakdowns and analyses to offer a more comprehensive view of the models’ capabilities on the Bongard OpenWorld dataset.
A.6.1 Componential Analysis Results by Commonsense Category
Table A.3 presents a detailed breakdown of the overall accuracy achieved by GPT-4o and Gemini 2.0 on the Bongard OpenWorld test set using the Componential Analysis paradigm. The results are categorized by Commonsense ID, providing insights into model performance across different types of commonsense reasoning challenges. This granular view allows for a more nuanced understanding of the models’ strengths and weaknesses in handling various aspects of commonsense knowledge.
ID Concept Category GPT 4o Gemini 2.0
0 Anything else 92.88 94.23
1 Human-Object Interaction (HOI) 86.67 92.86
2 Taste / Nutrition / Food 100.00 85.71
3 Color / Material / Shape 88.89 91.67
4 Functionality / Status / Affordance 100.00 100.00
5 And / Or / Not 90.00 80.00
6 Factual Knowledge 90.00 90.00
7 Meta Class 100.00 100.00
8 Relationship 100.00 100.00
9 Unusual Observations 92.31 92.31
Table A.3: Overall accuracy of GPT-4o and Gemini 2.0 on the Bongard OpenWorld test set using Componential Analysis, broken down by Commonsense ID.
A.6.2 Semantic Similarity in Deductive Rule Learning
Table A.4 presents the mean and standard deviation of the semantic similarity scores between the descriptions of the query images and the rules derived by the models during the Deductive Rule Learning experiments. These scores provide a measure of how well the extracted rules capture the semantic content of the images and the underlying Bongard problem. Analyzing these similarities helps to understand the relationship between the models’ ability to articulate rules and their overall performance on the task.
Gemini 2.0 GPT-4o
Category Mean Std Mean Std
Positive 0.915 0.02 0.9022 0.02
Negative 0.868 0.02 0.8659 0.02
Table A.4: Semantic Similarity between query image descriptions and identified Rules derived from Deductive Rule Learning
A.6.3 Impact of Chain-of-Thought Prompting
This section investigates the influence of Chain-of-Thought (CoT) prompting on VLM performance in the Direct Visual Rule Learning paradigm. CoT prompting, which encourages models to generate intermediate reasoning steps, has demonstrated significant improvements in complex reasoning tasks Wei et al. (2022b). We explore whether this approach is similarly beneficial in the visual domain, particularly for Bongard problems that demand a combination of perception and reasoning.
To assess the impact of CoT, we compare two prompting strategies: one that explicitly guides the model through a chain-of-thought by requesting a structured output (analysis, rule, query image description, and classification), and another that simply asks for the final classification.
Our analysis reveals a striking difference in performance between the two strategies. Without CoT prompting, the accuracy on the Direct Visual Rule Learning task drops significantly, from 80% to 61.60%. This decline is particularly pronounced for negative samples, where the accuracy falls from 66.40% to 39.20%, suggesting that the model struggles to identify instances where the rule does not apply when not explicitly prompted for a detailed reasoning process. This finding underscores the crucial role of CoT in guiding the model’s reasoning and enabling it to effectively navigate the complexities of visual reasoning tasks. These results are summarized in Table A.5.
Model Accuracy (%)
Overall neg pos
Without CoT 61.60 39.20 84.00
With CoT 80 66.40 93.60
Table A.5: Impact of CoT Prompting on Direct Visual Rule Learning when evaluated using GPT-4o model
The substantial impact of CoT observed in our experiments highlights its importance as a technique for eliciting more robust and reliable visual reasoning in VLMs. This suggests that incorporating CoT-like mechanisms into the model architecture or training process could lead to significant improvements in performance on complex visual tasks. Furthermore, our findings suggest that future research should explore the application of CoT prompting to other visual reasoning tasks and investigate the optimal strategies for integrating CoT into different VLM architectures.
A.6.4 Error Analysis Examples
This section provides detailed examples of Bongard OpenWorld test cases where both GPT-4o and Gemini 2.0 made incorrect classifications using the Componential Analysis paradigm. These examples illustrate common challenges faced by the models, offering insights into their limitations in visual reasoning and commonsense understanding. The analysis focuses on cases where both models failed, highlighting shared weaknesses. The “Reason” column explains the likely cause of misclassification based on the GPT-4o output (longer analysis, short rule, and test image description). These captions are reproduced from Wu et al. (2024).
Table A.6: Examples of Bongard OpenWorld Test Cases with Captions (Reproduced from Wu et al. (2024)) and Reasoning for Incorrect Classifications for images classified incorrectly by componential analysis by both by GPT-4o and Gemini 2.0. Human error analysis is based on the GPT-4o textual output: a longer analysis, short rule and the test image description.
SL NO Test ID Caption Reason
1 0021_neg_0 Cars on the city streets at night Weak reasoning (by similarity): a specific object in the rule (vehicles) is missing in the test image (impressionist painting), although deducible.
2 0014_neg_0 A person playing a guitar. Rule misses specific object (guitar) while being too general. The object is present in longer analysis.
3 0033_neg_0 A bicycle is placed in the corner Rule and test image analysis miss a specific property (in a corner / standing against a wall with ground visible). The test image is a collage of drawings on a wooden background.
4 0037_neg_0 The girl has long and thin braids on her head. Rule and test image analysis misses a specific property (specifically long and thin braids), being too general.
5 0076_pos_0 Various kinds of rings Rule misses a specific object (ring), being too general. Instead, rule focuses on a specific property (intricate design), which is missing in the test image and its description.
6 0076_neg_0 Various kinds of rings Rule misses a specific object (ring), being too general.
7 0082_neg_0 Live coral on the sea floor. Weak reasoning (by similarity): rule finds specific object (coral), but test image analysis does not.
8 0084_neg_0 A wooden fence surrounding a grassy field. Rule misses a specific object (grass), instead requiring greenery (grass or trees). The test image is a collage of natural images with greenery, no grass or trees are directly visible.
9 0112_neg_0 A wooden floor in the living room. Rule misses two specific objects (living room and a wooden floor), while focusing on just one object (wooden) and generalizing too much (indoor).
10 0117_neg_0 Colorful ribbons. Rule misses a specific object (colorful ribbons), generalizing too much.
11 0122_neg_0 A satellite view of Earth. Rule misses a specific viewpoint (top-down satellite view), using a too general (aerial view) instead.
12 0136_pos_0 Spectator seats view in the stadium. Weak reasoning: rule says “often active sports or spectators”, none of which are present in the test image or description. This lack is used as grounds for misclassification.
13 0213_neg_0 Checkerboard pattern fabrics Rule misses a specific object (fabric), the specific versions of which are present in analysis.
14 0234_neg_0 A beautiful stone sculpture Rule misses the correct property (tall), using instead another (prominent) of an obelisk. Test image features a prominent, but not a tall obelisk.
15 0247_pos_0 Small river filled with reeds Rule misses a specific object (reeds) present in analysis in a longer list, while prohibiting wrong objects (industrial presence), which are distantly present in test image.